# Quantum Algorithms for Matrix Operations Based on Unitary Transformations and Ancillary State Measurements

### v1

@@ Abstract
Matrix operations are of great significance in quantum computing, which manipulate quantum states in information processing. This paper presents quantum algorithms for several important matrix operations. By leveraging multi-qubit Toffoli gates and basic single-qubit operations, these algorithms efficiently carry out matrix row addition, row swapping, trace calculation and transpose. By using the ancillary measurement techniques to eliminate redundant information, these algorithms achieve streamlined and efficient computations, and demonstrate excellent performance with the running time increasing logarithmically as the matrix dimension grows, ensuring scalability. The success probability depends on the matrix dimensions for the trace calculation, and on the matrix elements for row addition. Interestingly, the success probability is a constant for matrix row swapping and transpose, highlighting the reliability and efficiency.
@@ 1 Introduction
Matrix operations are indispensable tools in quantum algorithms, which play key roles in quantum state manipulations in quantum system evolutions, quantum information processing and quantum computing. The design of quantum algorithms and the implementation of quantum circuits often rely heavily on matrix operations. Both Shor’s algorithm and Grover’s search algorithm optimize their computational processes through matrix operations. The HHL algorithm also relies fundamentally on matrix operations and has found numerous applications in matrix-based protocols Wang ; Tak ; Berry , providing new directions and momentum for the study of matrix operations and inspiring the development of more quantum algorithms for matrix computations. Given the broad and significant applications of matrix operations across disciplines, using quantum computing to address algebraic problems related to matrices has become a prominent research focus.
In recent years, researchers have proposed several quantum algorithms for matrix operations. For example, the Trotter approximation formula THF has been utilized to simulate quantum circuits, addressing complex operations such as matrix addition, multiplication and Kronecker products. Quantum algorithms for computing vector inner products ZhaoZ ; SZ_2019 ; QZKW_arxive2022 ; ZQKW_arxiv2022; ZQKW_arxiv2023 ; ZQKW_arxiv2024 have also been developed. These algorithms leverage the superposition and entanglement of qubits and have potential applications in fields like machine learning and data processing. Based on a “sender-receiver” quantum computing model, in 2022 Wen-Tao Qi et al. QZKW_arxive2022 proposed a class of quantum algorithms for matrix operations, including matrix-vector multiplication, matrix-matrix multiplication, matrix addition, determinant computation and matrix inversion. In 2023 A.I. Zenchuk et al. ZQKW_arxiv2023 further developed the idea of using specific types of unitary transformations to implement linear algebra protocols according to multi-qubit Toffoli gates and the simplest single-qubit operations. Auxiliary measurements were employed to remove redundant information and extract desired results. Building on this work, in 2024 A.I. Zenchuk et al. ZQKW_arxiv2024 introduced a method to encode the $N$ rows of a matrix into pure states of $N$ independent quantum subsystems. By selecting appropriate unitary operators, the determinant can be computed efficiently.
@@ 2.1 Adding one row of a matrix to another row
For the quantum algorithm that adds one row of a matrix to another, the quantum circuit diagram is shown in FIG.1. This algorithm requires two pure states and four auxiliary states, and is divided into eight steps.
@@
The following steps give rise to the algorithm of adding the $(k+1)$th row of a matrix $A$ to the $(l+1)$th row with $10\leq l,k\leq N-$ 1.
@@
Since the control operator $W^{(1)}_{R_{2}B_{1}}$ has an $n$-qubit control register, it can be expressed by using $O(n)$ Toffoli gates KShV . Therefore, the circuit complexity for computing $|\Phi_{1}\rangle$ is $O(n)=O(\log\,N)$.
@@
To precisely extract the element located at the $k+1$th row of the second term in $|\Phi_{1}\rangle$, from the terms in $|\Phi_{1}\rangle$ labeled by $|0\rangle_{B_{1}}$ we only extract the terms where $R_{1}$ is in the state $|k\rangle_{R_{1}}$. To do so, we introduce a 1-qubit auxiliary $B_{2}$ in the state $|0\rangle_{B_{2}}$, and the following control operator,
@@
where $P_{R_{1}B_{1}}=|k\rangle_{R_{1}}|0\rangle_{B_{1}}\;{{}_{R_{1}}}\langle k|{{}_{B_{1}}}\langle 0|$ is a projection operator.
@@
Step 4: Use SWAP gate to exchange the states of $R_{1}$ and $R_{2}$ marked by $|1\rangle_{B_{2}}$.
To add the $(k+1)$th row to the $(l+1)$th row, we apply a SWAP gate operation to the terms in $|\Phi_{2}\rangle$ labeled by $|1\rangle_{B_{2}}$, which swaps the states of $R_{1}$ and $R_{2}$. This operation is given by the following gate,
@@
Note that the $SWAPs$ in the control operator $W^{(3)}_{R_{1}R_{2}B_{2}}$ have common single control qubit and are related to different pairs of qubits. Therefore, they can be applied simultaneously. Consequently, the depth of this operator is $O(1)$ KShV .
@@
The computational depth of the control operator $W^{(4)}_{B_{1}B_{2}B_{3}}$ is $O(2)$.
@@
The computational depth of the control operator $W^{(5)}_{B_{1}B_{2}}$ is $2$.
Step 7: Relabel the useful and redundant information.
Introduce an auxiliary qubit $B_{4}$ in state $|0\rangle_{B_{4}}$, along with the projection operator $P_{B_{1}B_{2}B_{3}}=|000\rangle_{B_{1}B_{2}B_{3}}\;{}_{B_{1}B_{2}B_{3}}\langle 000|$. Applying the control operator
$W^{(6)}_{B_{1}B_{2}B_{3}B_{4}}=P_{B_{1}B_{2}B_{3}}\otimes\sigma^{(x)}_{B_{4}}+(I_{B_{1}B_{2}B_{3}}-P_{B_{1}B_{2}B_{3}})\otimes I_{B_{4}}$
to $|\Phi_{5}\rangle|0\rangle_{B_{4}}$, we obtain
$|\Phi_{6}\rangle$ $=$ $W^{(6)}_{B_{1}B_{2}B_{3}B_{4}}|\Phi_{5}\rangle|0\rangle_{B_{4}}$
$+\sum_{j=0}^{M-1}a_{kj}|l\rangle_{R_{1}}|j\rangle_{C_{1}}\Big{)}|k\rangle_{R_{2}}|0\rangle_{B_{1}}|0\rangle_{B_{2}}|0\rangle_{B_{3}}|1\rangle_{B_{4}}$
$+|g_{2}\rangle_{R_{1}C_{1}R_{2}B_{1}B_{2}B_{3}}|0\rangle_{B_{4}}.$
The useful information terms are now labeled by $|1\rangle_{B_{4}}$, while the useless information terms are labeled by $|0\rangle_{B_{4}}$.
The computational depth of the control operator $W^{(6)}_{R_{2}B_{1}B_{2}B_{3}B_{4}}$ is $O(3)$.
Step 8: Measure the auxiliary state $B_{4}$ using $|1\rangle_{B_{4}}$.
By measuring the auxiliary $B_{4}$ in the state $|1\rangle_{B_{4}}$, we get
$|\Phi_{7}\rangle$ $=$ $|\Psi_{out}\rangle|k\rangle_{R_{2}}|000\rangle_{B_{1}B_{2}B_{3}},$ (8)
@@
$|\Psi_{out}\rangle=\frac{1}{G}\Big{(}\sum_{i\neq l}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}\\ +\sum_{j=0}^{M-1}(a_{lj}+a_{kj})|l\rangle_{R_{1}}|j\rangle_{C_{1}}\Big{)}$
@@
The overall computational depth of the algorithm is $O(n)$, determined by $W^{(1)}_{R_{2}B_{1}}$ and $W^{(2)}_{R_{1}B_{1}B_{2}}$.
@@ 2.2 Swapping two rows of a matrix
For the quantum algorithm that swaps two rows of a matrix, the quantum circuit diagram is shown in FIG.2. The algorithm requires two pure states and four auxiliaries, and is divided into eight steps to swap the $(k+1)$th and $(l+1)$th rows of matrix $A$ ($10\leq l,k\leq N-$ 1).
@@
where $P_{R_{2}C_{2}}=|l\rangle_{R_{2}}|k\rangle_{C_{2}}{{}_{R_{2}}}\langle l|{{}_{C_{2}}}\langle k|$ is the projection operator. Applying $W^{(1)}_{R_{2}C_{2}B_{1}}$ to $|\Phi_{0}\rangle|0\rangle_{B_{1}}$, we get
@@
The circuit complexity for computing $|\Phi_{1}\rangle$ is $O(n)=O(\log\,N)$.
@@
where $P_{R_{1}R_{2}}=|k\rangle_{R_{1}}|l\rangle_{R_{2}}\;{{}_{R_{1}}}\langle k|{{}_{R_{2}}}\langle l|$ and $P_{R_{1}C_{2}}=|l\rangle_{R_{1}}|k\rangle_{C_{2}}\;{{}_{R_{1}}}\langle l|{{}_{C_{2}}}\langle k|$ are projection operators. $\sigma^{(x)}_{{B_{2}}^{i}}$ and $I_{{B_{2}}^{i}}$ are the operators acting on the $i$-th qubit of the auxiliary $B_{2}$ $(i=1,2)$.
@@
The computational depth of the control operator $W^{(2)}_{R_{1}R_{2}C_{2}B_{2}}$ is $O(2n)=O(\log\,N)$.
Step 4: Use SWAP gates to swap the states of $R_{1}$ and $R_{2}$, as well as the states of $R_{1}$ and $C_{2}$.
To swap the $(k+1)$th and $(l+1)$th rows of the matrix, we apply SWAP gate to the terms in $|\Phi_{2}\rangle$ labeled by $|10\rangle_{B_{2}}$ and $|01\rangle_{B_{2}}$, so as to swap the states of $R_{1}$ and $R_{2}$, as well as the states of $R_{1}$ and $C_{2}$. Denote
@@
The circuit complexity for computing $W^{(3)}_{R_{1}R_{2}B_{2}}$ is $O(1)$.
@@
The control operator $W^{(4)}_{B_{1}B_{2}B_{3}}$ has a computational depth of $3$.
@@
At this point, we have successfully swaped the elements of the $(k+1)$-th row of matrix $A$ with the corresponding elements of the $(l+1)$-th row. The control operator $W^{(5)}_{B_{1}B_{2}}$ has a computational depth of $3$.
Step 7: Re-label the useful information terms and useless information terms.
Since new useless information terms are generated in the previous operation, and the useful information is stored in the first term of $|\Phi_{5}\rangle$, we need to re-mark the useful and useless information terms. To do so, we introduce an auxiliary qubit $B_{4}$ in state $|0\rangle_{B_{4}}$ and a projection operator $P_{B_{1}B_{2}B_{3}}=|0001\rangle_{B_{1}B_{2}B_{3}}\;{{}_{B_{1}B_{2}B_{3}}}\langle 0001|.$ We apply the following control operator
$W^{(6)}_{B_{1}B_{2}B_{3}B_{4}}=P_{B_{1}B_{2}B_{3}}\otimes\sigma^{(x)}_{B_{4}}+(I_{B_{1}B_{2}B_{3}}-P_{B_{1}B_{2}B_{3}})\otimes I_{B_{4}}$
to state $|\Phi_{5}\rangle|0\rangle_{B_{4}}$, which gives rise to
$|\Phi_{6}\rangle$ $=$ $W^{(6)}_{B_{1}B_{2}B_{3}B_{4}}|\Phi_{5}\rangle|0\rangle_{B_{4}}$
$+\sum_{j=0}^{M-1}a_{lj}|k\rangle_{R_{1}}|j\rangle_{C_{1}}+\sum_{j=0}^{M-1}a_{kj}|l\rangle_{R_{1}}|j\rangle_{C_{1}}\Big{)}|l\rangle_{R_{2}}|k\rangle_{C_{2}}$
$|0\rangle_{B_{1}}|00\rangle_{B_{2}}|1\rangle_{B_{3}}|1\rangle_{B_{4}}+|g_{3}\rangle_{R_{1}C_{1}R_{2}C_{2}B_{1}B_{2}B_{3}}|0\rangle_{B_{4}},$
where the useful information terms are marked by $|1\rangle_{B_{4}}$, and the useless information terms are marked by $|0\rangle_{B_{4}}$.
The control operator $W^{(6)}_{B_{1}B_{2}B_{3}B_{4}}$ has a computational depth of $4$.
Step 8: Measure the auxiliary $B_{4}$ in the state $|1\rangle_{B_{4}}$
We measure the auxiliary $B_{4}$ in the state $|1\rangle_{B_{4}}$ to remove the useless information terms, and obtain
$|\Phi_{7}\rangle=|\Psi_{out}\rangle|l\rangle_{R_{2}}|k\rangle_{C_{2}}|0001\rangle_{B_{1}B_{2}B_{3}},$ (16)
@@
Clearly, the success probability of the measurement is $\frac{1}{24}$ and the overall depth of the algorithm is $O(n)$.
3 Matrix Trace Calculation and Matrix Transposition
@@ 3.1 Matrix Trace Calculation
The quantum circuit for calculating the trace of the matrix $S$ is shown in FIG.3. Assume that a matrix $S$=$\{a_{ik}\}$ is of size $N\times N$ with $N=2^{n}$. This algorithm requires 1 pure state and 3 auxiliary qubits, and is divided into the following 6 steps.
@@
where $P^{(m)}_{j}=|m_{j}\rangle_{R}|m_{j}\rangle_{C}\;{{}_{R}}\langle m_{j}|{{}_{C}}\langle m_{j}|$, $m=0,1$, $j=1,2,...,n$, is the projection operator acting on the $j$-th qubit of registers $R$ and $C$, $\sigma^{(x)}_{A^{(j)}}$ is the Pauli matrix acting on the $j$-th qubit of auxiliary $A$, $I_{A^{(j)}}$ is the identity operator acting on the $j$-th qubit of auxiliary $A$.
@@
The computational depth for obtaining $|\Phi_{2}\rangle$ is $O(n)=O(\log\,N)$.
@@
Since new useless information terms were generated in the previous step and the useful information is stored in the first term of $|\Phi_{3}\rangle$, we need to re-mark the useful and useless information terms. To do this, we introduce an auxiliary qubit $B_{2}$ in state $|0\rangle_{B_{2}}$, as well as the projection operator $P_{RCAB_{1}}=|0\rangle_{R}|0\rangle_{C}|0\rangle_{A}|1\rangle_{B_{1}}\;{{}_{R}}\langle 0|{{}_{C}}\langle 0|_{A}\langle 0|_{B_{1}}\langle 1|.$ We apply the following control operator
@@
The control operator $W^{(4)}_{S_{1}S_{2}AB_{1}B_{2}}$ has $13n+$ 1 control qubits, and can be represented using $O(3n)=O(n)$ Toffoli gates. So it takes $O(n)=O(\log\,N)$ steps to compute $|\Phi_{5}\rangle$.
@@
The trace is obtained based on the measurement probability $\frac{|\sum_{i=0}^{N-1}a_{ii}\rangle|^{2}}{2^{3n}}$.
The overall depth of the algorithm given by the operators $W^{(2)}_{AB_{1}}$ and $W^{(4)}_{RCAB_{1}B_{2}}$ is $O(n)=O(logN)$.
3.2 Matrix Transposition
The quantum circuit for matrix transposition is shown in FIG.4. Assume that a matrix $S$=$\{a_{ik}\}$ is of size $N\times M$ with $N=2^{n}$ and $M=2^{m}$. The algorithm requires 2 pure states and 1 auxiliary qubit, and is divided into 3 steps.
Figure 4: Quantum Circuit for Matrix Transposition.

### v2

# Quantum Algorithms for Matrix Operations of Row Addition, Row Swapping, Trace Calculation and Transpose

@@ Abstract
Quantum algorithms of matrix operations are of great significance in many fields in science and technology. In this paper, by leveraging multi-qubit Toffoli gates and basic single-qubit operations, the quantum algorithms of matrix operations of row addition, row swapping, trace calculation and transpose are obtained. In particular, the complexities of these quantum algorithms are presented, too.
@@ 1 Introduction
Matrix operations serve as the core computational power support for quantum algorithms. The efficient quantum implementation of its basic operations (including row transformation, transposition, and trace calculation, etc.) holds significant theoretical and practical value. In quantum chemistry simulations MYM , the elementary row transformation of matrices can rapidly simplify the Hamiltonian matrix, reducing the time complexity of solving molecular energy spectra from the polynomial level of classical algorithms to the logarithmic level. Matrix Transpose operation enables the efficient analysis of molecular symmetry, while the calculation of the matrix trace can directly yield key physical quantities such as the number of electrons. In quantum cryptography KSG , elementary row transformation can optimize the encoding of quantum states to enhance security. Matrix Transpose can achieve efficient information routing in quantum communication networks, and the matrix trace can be used for the integrity detection of quantum states. In the field of quantum machine learning PA , elementary row transformation can accelerate the feature extraction of Quantum Support Vector Machines (QSVMs). Matrix Transpose can improve the backpropagation efficiency of quantum neural networks, and the matrix trace operation can quickly calculate kernel functions to evaluate model performance. Given the wide-ranging influence of matrix operations in interdisciplinary fields, exploring efficient quantum computing methods to solve matrix-related algebraic problems has become a current research hotspot.
In recent years, researchers have proposed several quantum algorithms for matrix operations. For example, the Trotter approximation formula THF has been utilized to simulate quantum circuits, addressing complex operations such as matrix addition, multiplication and Kronecker products. Quantum algorithms for computing vector inner products ZhaoZ ; SZ_2019 ; QZKW_arxive2022 ; ZQKW_arxiv2023 ; ZQKW_arxiv2024 have also been developed. These algorithms leverage the superposition and entanglement of qubits and have potential applications in fields like machine learning and data processing. Based on a “sender-receiver” quantum computing model, in 2022 Wen-Tao Qi et al. QZKW_arxive2022 proposed a class of quantum algorithms for matrix operations, including matrix-vector multiplication, matrix-matrix multiplication, matrix addition, determinant computation and matrix inversion. In 2023 A.I. Zenchuk et al. ZQKW_arxiv2023 further developed the idea of using specific types of unitary transformations to implement linear algebra protocols according to multi-qubit Toffoli gates and the simplest single-qubit operations. Auxiliary measurements were employed to remove redundant information and extract desired results. Building on this work, in 2024 they ZQKW_arxiv2024 also introduced a method to encode the $N$ rows of a matrix into pure states of $N$ independent quantum subsystems. By selecting appropriate unitary operators, the determinant can be computed efficiently. Furthermore, in 2025 they ZQW_arxiv2025 discussed the method of encording unnormalized matrices, and the system encoding the matrix was extended adding one more probability amplitude that satisfy the normalization condition.
@@
This paper focuses on the core algorithm layer of quantum computers and aims to design a set of quantum circuit schemes for implementing basic matrix operations. The research is carried out under the following premise: the input state, which is the quantum state encoding matrix elements, is successfully prepared. On this basis, we systematically construct a quantum circuit architecture capable of performing basic matrix operations such as row transformation transposition, and trace calculation.
This approach is of great value in the current development of quantum computing. On the one hand, quantum state preparation and quantum state measurement, as crucial aspects in quantum computing, have attracted much attention on conducting specialized studies. Transforming classical matrices into quantum pure states is a fundamental issue that needs to be addressed in the field of quantum state preparation. Moreover, converting the results of quantum state measurements into classical data interpretable by humans makes quantum readout the core part of quantum state measurement research. On the other hand, the algorithm designed in this paper fully considers the compatibility with quantum state preparation and quantum state measurement schemes, laying a solid foundation for the rapid deployment and application of quantum computers after they become practically implementable in the future.
@@ 2.1 Adding one row of a matrix to another row
For the quantum algorithm that adds one row of a matrix to another, the quantum circuit diagram is shown in FIG.1. This algorithm requires two pure states and four auxiliary states, and is divided into seven steps.
@@
The following steps give rise to the algorithm of adding the $(k+1)$th row of a matrix $A$ to the $(l+1)$th row with $0\leq l,k\leq N-1$.
@@
Kitaev A. Y. et. al KShV represented an important result: A control operator with $n$ control qubits can be represented by $O(n)$ Toffoli gates. Since the above control operator $W^{(1)}_{R_{2}B_{1}}$ has an $n$-qubit control register, it can be expressed by using $O(n)$ Toffoli gates. Therefore, the circuit complexity for computing $|\Phi_{1}\rangle$ is $O(n)=O(\log\,N)$.
@@
To precisely extract the element located at the $(k+1)$th row of the second term in $|\Phi_{1}\rangle$, from the terms in $|\Phi_{1}\rangle$ labeled by $|0\rangle_{B_{1}}$ we only extract the terms where $R_{1}$ is in the state $|k\rangle_{R_{1}}$. To do so, we introduce a 1-qubit auxiliary $B_{2}$ in the state $|0\rangle_{B_{2}}$, and the following control operator,
@@
[…]
@@
The control operator $W^{(2)}_{R_{1}B_{1}B_{2}}$ has $n+1$ control qubits, so it can be represented in terms of $O(n)$ Toffoli gates and the circuit complexity for computing $|\Phi_{2}\rangle$ is $O(n)=O(\log\,N)$.
Step 4: Use C-SWAP gate to exchange the states of $R_{1}$ and $R_{2}$ marked by $|1\rangle_{B_{2}}$.
To add the $(k+1)$th row to the $(l+1)$th row, we apply a C-SWAP gate operation to the terms in $|\Phi_{2}\rangle$ labeled by $|1\rangle_{B_{2}}$, which swaps the states of $R_{1}$ and $R_{2}$. This operation is given by the following gate,
@@
Note that the $SWAPs$ in the control operator $W^{(3)}_{R_{1}R_{2}B_{2}}$ have common single control qubit and are related to different pairs of qubits. Therefore, they can be applied simultaneously.
@@
Since the control operator $W^{(4)}_{B_{1}B_{2}B_{3}}$ has two control qubit, its computational complexity is $O(1)$.
@@
Since only two Hadamard gates are used here, the circuit complexity is $O(1)$.
Step 7: Measure the auxiliary state $B_{1}B_{2}B_{3}$ using $|000\rangle_{B_{1}B_{2}B_{3}}$.
By measuring the auxiliary $B_{1}B_{2}B_{3}$ in the state $|000\rangle_{B_{1}B_{2}B_{3}}$, we get
$|\Phi_{6}\rangle$ $=$ $|\Psi_{out}\rangle|k\rangle_{R_{2}},$ (7)
@@
$|\Psi_{out}\rangle$ $=$ $\frac{1}{G}\Big{(}\sum_{i\neq l}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}$
$+\sum_{j=0}^{M-1}(a_{lj}+a_{kj})|l\rangle_{R_{1}}|j\rangle_{C_{1}}\Big{)}$
@@
The overall computational complexity of the algorithm is $O(n)$, mainly determined by $W^{(1)}_{R_{2}B_{1}}$ and $W^{(2)}_{R_{1}B_{1}B_{2}}$.
@@ 2.2 Swapping two rows of a matrix
For the quantum algorithm that swaps two rows of a matrix, the quantum circuit diagram is shown in FIG.2. The algorithm requires two pure states and four auxiliaries, and is divided into seven steps to swap the $(k+1)$th and $(l+1)$th rows of matrix AA ($0\leq l,k\leq N-1$).
@@
[…]
@@
Since the control operator $W^{(1)}_{R_{2}C_{2}B_{1}}$ has $n$ control qubit, its computational complexity is $O(n)=O(\log\,N)$.
@@
[…]
@@
Since the control operator $W^{(2)}_{R_{1}R_{2}C_{2}B_{2}}$ has $2n$ control qubit, its computational complexity is $O(n)=O(\log\,N)$.
Step 4: Use C-SWAP gates to swap the states of $R_{1}$ and $R_{2}$, as well as the states of $R_{1}$ and $C_{2}$.
To swap the $(k+1)$th and $(l+1)$th rows of the matrix, we apply C-SWAP gate to the terms in $|\Phi_{2}\rangle$ labeled by $|10\rangle_{B_{2}}$ and $|01\rangle_{B_{2}}$, so as to swap the states of $R_{1}$ and $R_{2}$, as well as the states of $R_{1}$ and $C_{2}$. Denote
@@
Note that the $SWAPs$ in the control operator $W^{(3)}_{R_{1}R_{2}B_{2}}$ have common single control qubit and are related to different pairs of qubits. Therefore, they can be applied simultaneously.
@@
Since the control operator $W^{(4)}_{B_{1}B_{2}B_{3}}$ has one control qubit, its computational complexity is $O(1)$.
@@
At this point, we have successfully swaped the elements of the $(k+1)$-th row of matrix $A$ with the corresponding elements of the $(l+1)$-th row.
Since only three Hadamard gates are used here, the circuit complexity is $O(1)$.
Step 7: Measure the auxiliary $B_{1}B_{2}B_{3}$ in the state $|0001\rangle_{B_{1}B_{2}B_{3}}$
We measure the auxiliary $B_{1}B_{2}B_{3}$ in the state $|0001\rangle_{B_{1}B_{2}B_{3}}$ to remove the useless information terms, and obtain
$|\Phi_{6}\rangle=|\Psi_{out}\rangle|l\rangle_{R_{2}}|k\rangle_{C_{2}},$ (15)
@@
The success probability is $\frac{1}{24}$. This probability is independent of the matrix dimension.
The overall computational complexity of the algorithm is $O(n)$, mainly determined by $W^{(1)}_{R_{2}C_{2}B_{1}}$ and $W^{(2)}_{R_{1}R_{2}C_{2}B_{2}}$.
For specific examples of this algorithm, please refer to Appendix 1.
3 Matrix Trace Calculation and Transpose
@@ 3.1 Matrix Trace Calculation
[…]
@@
[…]
@@
The control operator $W^{(1)}_{RCA}$ has $2n$ control qubits, so the circuit complexity for computing $|\Phi_{1}\rangle$ is $O(n)$.
@@
Since the control operator $W^{(2)}_{AB_{1}}$ has $n$ control qubits, its computational complexity is $O(n)=O(\log\,N)$.
@@
Since $3n$ Hadamard gates are used here, the circuit complexity is $O(n)$.
@@
[…]
@@
Since the control operator $W^{(4)}_{RCAB_{1}B_{2}}$ has $3n+1$ control qubit, its computational complexity is $O(n)$.
@@
The trace is obtained based on the measurement probability $\frac{|\sum_{i=0}^{N-1}a_{ii}|^{2}}{2^{3n}}$. This probability is related of the matrix dimension.
The overall computational complexity of the algorithm is $O(n)$, mainly determined by $W^{(1)}_{RCA}$, $W^{(2)}_{AB_{1}}$, $W^{(3)}_{RCA}$ and $W^{(4)}_{RCAB_{1}B_{2}}$.
3.2 Matrix Transpose
[…]
Figure 4: Quantum Circuit for Matrix Transpose.